OpenStudy (anonymous):

A refrigrator is sold for Rs.11475 after giving succesive discounts of 15% and 10% respectively.Find the marked price of the refrigrator.

OpenStudy (anonymous):

I ONLY HAVE 2 MINUTES LEFT PPPPPLLLLLLLSSSSSSSSS

OpenStudy (anonymous):

why

OpenStudy (anonymous):

BS THERE IS A TEST IN MY COACHING

OpenStudy (anonymous):

Cum onnn

OpenStudy (anonymous):

10975 approx not sure

OpenStudy (anonymous):

the answer should be 15000

OpenStudy (anonymous):

ok so u know the answer

OpenStudy (anonymous):

but i need the solution

OpenStudy (anonymous):

14404.875

OpenStudy (anonymous):

ok can u show me the solution.

OpenStudy (anonymous):

i was already hving the ans. written in my book

OpenStudy (anonymous):

give me the solution

OpenStudy (anonymous):

11475/(1-0.15)=13500 13500/(1-0.1)=15000

OpenStudy (anonymous):

but the discounts are 15 and 20 percent

OpenStudy (anonymous):

oho u said 15 and 10 % gary

OpenStudy (anonymous):

A refrigrator is sold for Rs.11475 after giving succesive discounts of 15% and 10% respectively.Find the marked price of the refrigrator.

OpenStudy (anonymous):

sorry but still 15 % not equals to 0.15

OpenStudy (anonymous):

15% = 15/100 = 0.15

OpenStudy (anonymous):

thanks everyone

OpenStudy (anonymous):

my pleasure

OpenStudy (anonymous):

ok bye wish me luck for the test

OpenStudy (anonymous):

They said the discounts were 15% and 10% respectively, so we have to undo the 10% discount first, and the 15% discount second. 11475/(1-0.1) = 11475/0.9 = 12750 12750/(1-0.15) = 12750/0.85 = 15000 Taking 90% of 85% should equal 85% of 90%, so going backward in either order should get you to the same place. But this is just how I read the question.